HUBBLE SPACE TELESCOPE DAILY REPORT #1377 PERIOD COVERED: 0000Z (UTC) 04/25/95 - 0000Z (UTC) 04/26/95 Daily Status Report as of 116/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5560 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 1) The WF/PC-2 was used to make several internal bias, K-Spot, and flat calibration observations to monitor the CCD camera electronics, the optical alignment, and to check for contamination. The observations were completed as planned, and no anomalies were noted. 1.2 Completed WF/PC-2 5496 (M-3: An Ideal Laboratory for Testing Stellar Evolution and Dynamics) The WF/PC-2 was used to make observations of globular cluster NGC-5272 (aka, M-3) to identify the evolved stars, and to determine their luminosity function. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Two Sets of WF/PC-2 5446 (The Nuclei of Nearby S0 and Spiral Galaxies: A PC Snapshot Imaging) The WF/PC-2 was used to make observations of the core region of galaxy UGC-8085 to provide statistics on the nuclear cusps, stellar condensations, and inner stellar disks and dust lanes. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 6248 (HST Observations of Nova Scorpii) The WF/PC-2 was used to make observations of the superluminal transient X-Ray source GRO-J1655 (aka, X-Ray Nova Scorpii 1994) following the outburst which started on March 20, 1995. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 4 There was a guide star acquisition failure at 116/0858Z (HSTAR 5193). No star was found, and the system did not fall back into coarse track. Subsequent reacquisitions were also affected. Analysis of the anomaly is in progress. Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled : 12 Successful : 12 2.3 Significant HSTARs: None 2.4 Operations Notes: During one of the two dumps of ESTR #2, there were two bit slips. The Battery #3 capacity test is underway. Discharge was started at 115/1340Z and terminated at 116/0550Z. Battery #3 was connected to SA Section 3 at 116/0747Z. Charging was begun at 116/0753Z. It is estimated that the six battery configuration will be achieved at approximately 117/1950Z. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-